OpenStudy (anonymous):

will medal Solve for the variable r in this equation: s = 2πrh

OpenStudy (smilie2015):

as pi ,2 and h are multiplying here so divide both sides with pi,2,h

OpenStudy (smilie2015):

\[s/2\pi h = r\]

OpenStudy (lovelovelovehelp):

r = s/ 2πh